HOW MANY CUBIC INCHES OF MATERIAL REQUIRED FOR SOLID RUBBER BALL WHICh DIAMETER OF 3 INCH?
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

HOW MANY CUBIC INCHES OF MATERIAL REQUIRED FOR SOLID RUBBER BALL WHICh DIAMETER OF 3 INCH?
--------------------
V = (4/3)pi*r^3
----
Your Problem:
V = (4/3)pi*(3/2)^3
---
V = 14.137.. cubic inches
==============================
Cheers,
Stan H.

Answer by checkley77(12844)

V=4/3PIR^3
V=4/3*3.14*3/2^3
V=4.18666*27/8
V=14.13 IN^3 ANS.